40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Feb 10 22:21:55 2014, manasa, Update, Green Locking, X and Y arm green tuned 
    Reply  Tue Feb 11 21:22:02 2014, manasa, Update, Green Locking, ALS X and Y arm restored 
       Reply  Tue Feb 11 22:17:06 2014, Koji, Update, Green Locking, ALS X and Y arm restored 
    Reply  Thu Feb 13 16:48:33 2014, Steve, Update, Green Locking, X and Y arm green tuned details 
Message ID: 9624     Entry time: Tue Feb 11 21:22:02 2014     In reply to: 9621     Reply to this: 9625
Author: manasa 
Type: Update 
Category: Green Locking 
Subject: ALS X and Y arm restored 

The X and Y arms were locked successfully using ALS and the arms could be scanned and held to support IR resonance.

The same procedure as in elog 9219 was followed. In-loop noise was measured to be between 200-300 Hz rms for the lock.

ALS settings for the lock

X arm : FM 2, 3, 5, 6, 7, 8, 10  Gain = 11.0
Y arm : FM 2, 3, 5, 6, 7, 8, 10  Gain = 10.0

ELOG V3.1.3-